OpenStudy (anonymous):

A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue: Write two functions to model this situation, and graph those two functions on the same coordinate grid. During what year in the future will the park have approximately the same number of pine and oak trees? How many of each type of tree will there be at that time?

OpenStudy (anonymous):

@superhelp101 @hero please help

OpenStudy (paxpolaris):

number of pine trees \[= 800\cdot(1-0.05)^t=800\times0.95^t\] number of oaktrees \[= 50\cdot(1+0.15)^t=50\times1.15^t\]

OpenStudy (anonymous):

thanks i got it from here